# Emergent Hierarchical Reasoning in LLMs through Reinforcement Learning

### v2

@@ Abstract
Reinforcement Learning (RL) has proven highly effective at enhancing the complex reasoning abilities of Large Language Models (LLMs), yet underlying mechanisms driving this success remain largely opaque. Our analysis reveals that puzzling phenomena like “aha moments”, “length-scaling” and entropy dynamics are not disparate occurrences but hallmarks of an emergent reasoning hierarchy, akin to the separation of high-level strategic planning from low-level procedural execution in human cognition. We uncover a compelling two-phase dynamic: initially, a model is constrained by procedural correctness and must improve its low-level skills. The learning bottleneck then decisively shifts, with performance gains being driven by the exploration and mastery of high-level strategic planning. This insight exposes a core inefficiency in prevailing RL algorithms like GRPO, which apply optimization pressure agnostically and dilute the learning signal across all tokens. To address this, we propose HIerarchy-Aware Credit Assignment (HICRA), an algorithm that concentrates optimization efforts on high-impact planning tokens. HICRA significantly outperforms strong baselines, demonstrating that focusing on this strategic bottleneck is key to unlocking advanced reasoning. Furthermore, we validate semantic entropy as a superior compass for measuring strategic exploration over misleading metrics such as token-level entropy.
@@ 1 Introduction
In this work, we posit that these phenomena in effective RL processes are not disparate occurrences but are hallmarks of a single, coherent process: RL forges an emergent functional reasoning hierarchy within the LLM. We argue that the learning process dynamically shifts its focus, first building a reliable foundation of low-level procedural skills and then learning to direct these skills with high-level strategic planning. This emergent structure finds a compelling parallel in the cognitive architecture of the human brain (Murray et al., 2014; Zeraati et al., 2023; Huntenburg et al., 2018), which separates high-level, deliberate strategic planning (e.g., deciding which theorem to apply in the prefrontal cortex) from the rapid, subordinate execution of learned procedures (e.g., performing an algebraic manipulation).
This cognitive model provides a powerful framework for interpreting the learning dynamics we observe in LLMs. Motivated by this parallel, we propose a functional decomposition of model-generated solutions:
@@
Our analysis across various models reveals that the optimization pressure on the reasoning hierarchy evolves throughout training, driven by a dynamic shift in the learning frontiers. At the outset, the learning process is constrained by procedural correctness. A single calculation error can nullify an entire solution, providing a strong learning signal that forces the model to first master low-level execution tokens. Once the model achieves proficiency in these foundational skills, however, the learning bottleneck shifts to strategic planning. We find that exploring and mastering the use of planning tokens unlocks significant and sustained improvements in reasoning ability.
This two-phase mechanism explains ”aha moments” as the discovery and internalization of high-level strategic reasoning strategies, such as self-reflections. It also accounts for the ”length-scaling” effect, as employing more sophisticated strategies – involving thorough planning and logical backtracing – naturally elongates the reasoning trace with structured, strategic deliberation. Notably, it provides a unified perspective to understand the complex token entropy dynamics across different models, through the lens of high-impact planning tokens and gradually confident execution tokens.
@@
Our work makes three primary contributions:
  • We identify and empirically validate the emergent reasoning hierarchy in RL-tuned LLMs. We, for the first time, show that the enhanced reasoning is not a monolithic process but a two-phase dynamic: The learning frontier shifts from mastering low-level procedural execution to exploring high-level strategic planning. This core insight provides a unified explanation for opaque phenomena such as “aha moments”, “length-scaling” and the distinctive token entropy dynamics across different models.
  • To effectively address the strategic bottleneck, we introduce HICRA (Hierarchy-Aware Credit Assignment), an algorithm that concentrates optimization efforts on high-impact planning tokens. By moving beyond agnostic credit assignment, HICRA directly targets the primary driver of advanced reasoning (Section 3).
  • Through extensive experiments, we show that HICRA significantly outperforms strong baselines across multiple models and benchmarks through the lens of strategic exploration. In addition, we validate the use of semantic entropy as a reliable compass for measuring strategic exploration than conventional metrics like token-level entropy, which we show can be misleading.
@@ 2 The Emergent Reasoning Hierarchy
Our empirical analysis reveals that the enhancement of reasoning in LLMs through RL is not a monolithic process. Instead, it is driven by the emergence of a functional reasoning hierarchy, which forms in two overlapping phases. At the outset, the model’s learning is often constrained by the need for procedural reliability; it must first learn to execute low-level steps reliably. Once this foundation is built, the bottleneck shifts to strategic planning, where the exploration and mastery of high-level reasoning strategies become the primary engine for performance gains.
@@ 2.1 A Functional Proxy for the Reasoning Hierarchy
To analyze these dynamics, we must first distinguish between high-level planning and low-level tokens. This is inherently challenging, as a token’s function is not intrinsic but is defined by its role in a sequence.
We motivate our approach by drawing a parallel to human cognition. When a person reasons through a problem, we identify their strategic thinking by its function. A phrase like, “Let’s try a different approach,” functions as a high-level strategic maneuver that guides the problem-solving direction. In contrast, a phrase like, “so we add 5 to both sides,” is a low-level procedural step. Inspired by this functional distinction, we introduce Strategic Grams as a functional proxy to circumvent the difficulty of formally defining a “planning token”.
Strategic Grams (SGs) are defined as $n$-grams that function as a single semantic unit to guide the logical flow. We use n-grams because they capture the phrasal nature of strategic language (e.g., ”let’s consider the case”) which is lost at the single-token level. As Figure 2 shows, these SGs facilitate three main types of logical moves: (a) deduction, (b) branching, and (c) backtracing.
Given a collection of SGs, our classification heuristic is straightforward:
A token is classified as a strategic planning token if it is part of a Strategic Gram in the current context. All other tokens are classified as procedural execution tokens.
For simplicity, we use the term “execution tokens” to refer to all non-planning tokens. We note this is a slight simplification, as this category encompasses not only concrete calculations but also formatting and other procedural language.
We constructed the SG set using the following procedure:
@@
    Grams Collection: We generate a large corpus of successful solutions by sampling from the training model at different iterations, and then extract $n$-grams (where $n\in[3,5]$) to collect the model’s native mode of expression.
@@
    Grams Annotation: We use Gemini-2.5 with carefully designed prompts (see Appendix) to programmatically classify $n$-grams that function as strategic constructs, yielding a ”seed collection.” We then expand the seed set by manually reviewing any missing SGs from the rollout instances.
Using this heuristic and the resulting SG collection (listed in the Appendix), our empirical results in the following sections reveal consistent and insightful patterns in the model’s learning dynamics.
@@ 2.2 Emergence of the Reasoning Hierarchy
Building on our functional proxy for reasoning, we examine the learning dynamics of RL for LLM reasoning and finds an intriguing parallel with human-like hierarchical reasoning. Our empirical analysis – conducted consistently across different model families, Qwen2.5-7B (Yang et al., 2024), Qwen3-4B (Yang et al., 2025), LLama-3.1-8B (Grattafiori et al., 2024), MiMO-VL-7B (Xiaomi, 2025) – reveals that enhanced reasoning is not a monolithic process, but driven by an evolution of the learning frontiers.
The learning process exhibit two overlapping phases: it begins with a rapid consolidation of procedural reliability, conducive to the widespread low-level tokens. This is followed by a sustained period where the greatest potential for improvement shifts to the exploration of high-level strategic reasoning, which serves as the true engine of advanced performance.
2.2.1 Forging a Reliable Procedural Engine
The initial phase of RL training is dedicated to mastering the basics. The model must first build a reliable engine for performing calculations and other procedural steps. To observe this, we track two key metrics on the execution tokens:
  • Relative Perplexity: Perplexity, $P(x_{t}|x_{<t})$, measures model surprise. A lower value signifies higher confidence. We normalize the perplexity by its initial value to compare the rates of change in planning tokens and execution tokens.
  • Token-Level Entropy: The Shannon entropy of the policy’s next-token distribution, $H(\pi(\cdot|x_{<t}))$, measures its uncertainty. High entropy signals active exploration of the vocabulary at the position, while low entropy suggests confident exploitation of a known token.
@@ 2.2.1 Forging a Reliable Procedural Engine
We indeed find that the phase of procedural consolidation might be absent or shot in models with stronger capacity, as evident in MiMO-VL-Instruct and Qwen-4B-Instruct. We refer the reader to check the full analysis of training dynamics across six models in the appendix. This leads us to the belief that the primary driver of RL is indeed the exploration of strategic planning.
@@ 2.2.2 Steering the Skills with Strategic Planning
Once the model becomes procedurally reliable, its performance gains are primarily driven by its ability to explore and deploy a diverse set of high-level strategies. To track this shift, we analyze the planning tokens using two key metrics. We compute the Semantic Entropy of strategic grams – he Shannon Entropy of the frequency distribution of strategic grams – to quantify the diversity of the model’s high-level strategic plans. To isolate procedural variety, we also compute the conditional entropy of subsequent procedural n-grams given a preceding strategic gram. This second metric measures how varied is the subsequent procedural steps for a preceding strategic move.
@@
Explaining Puzzling Phenomena.
This emergent reasoning hierarchy provides a unified explanation for previously observed behaviors.
@@
Semantic Entropy: A Good Compass for Exploration.
The trends in Figure 3 also highlight a critical flaw in using aggregate token-level entropy (column 2) to track exploration.
@@
The aggregate token-level entropy is dominated by the vast majority of low-level execution tokens. As the model become confident in procedural steps, the entropy of these tokens naturally decreases, pulling the global average down.
Unluckily, the decrease in token-level entropy sometimes mislead practitioners into the conception of declined exploration. This is incorrect, however, as it contradicts the fact of increasing exploration in strategic plans (semantic entropy of planning tokens) and the improving reasoning performance. We refer interested readers to the appendix for a full analysis of training dynamics across six LLM and VLMs in the appendix.
Semantic entropy avoids this pitfall by directly measuring diversity at the semantic level of meaningful strategic units. Its trend accurately reflects the expansion of the model’s strategic playbook, making it a more reliable diagnostic tool for tracking genuine exploration and predicting sustained performance improvements. In Section 4.4, we showcase how semantic entropy avoids the flaws in token entropy and complements Pass@K metric with further benefits.
@@ 3 HICRA: Hierarchy-Aware Credit Assignment
Our empirical analysis reveals a fundamental insight into how Large Language Models (LLMs) learn to reason via Reinforcement Learning (RL): the learning process is characterized by a dynamic shift in its learning bottleneck. Initially, the model is constrained by procedural correctness, but as it masters these foundational skills, the frontier for performance improvement shifts to the exploration and mastery of high-level strategic planning.
@@
Formulation.
We introduce Hierarchy-Aware Credit Assignment (HICRA), an algorithm that builds upon the GRPO framework to allocate credit based on the reasoning hierarchy. In GRPO, given a query $\mathbf{q}$ from a dataset $\mathcal{D}$, the policy $\pi_{\theta}$ generates a set of $G$ output trajectories $\{\mathbf{o}_{1},\dots,\mathbf{o}_{G}\}$. The advantage for a token $o_{i,t}$ at timestep $t$ in trajectory $\mathbf{o}_{i}$ is the group-normalized reward:
$\hat{A}_{i,t}=R(\mathbf{q},\mathbf{o}_{i})-\frac{1}{G}\sum_{j=1}^{G}R(\mathbf{q},\mathbf{o}_{j})$
@@
Connection to Strategic Exploration.
The core mechanism of HICRA engineers more effective exploration by reshaping the policy update’s target distribution. A standard policy gradient (Williams, 2004) update nudges the policy $\pi_{\theta_{old}}$ toward an implicit target distribution $\pi^{\ast}$ defined by the advantage function described as follows (the derivation is included in the appendix):
@@
This anisotropic reshaping fosters a potent virtuous feedback loop: (a) the policy is incentivized to explore the subspace of strategic plans more thoroughly; (b) this leads to the faster discovery of effective reasoning patterns; and (c) when these strategies yield high rewards, the amplified advantage ensures they are strongly reinforced, cementing the model’s planning capabilities far more efficiently. We also validate the effects of HICRA in exploration through extensive experiments.

### v3

@@ Abstract
Reinforcement Learning (RL) has proven highly effective at enhancing the complex reasoning abilities of Large Language Models (LLMs), yet underlying mechanisms driving this success remain largely opaque. Our analysis reveals that puzzling phenomena like “aha moments”, “length-scaling” and entropy dynamics are not disparate occurrences but hallmarks of an emergent reasoning hierarchy, akin to the separation of high-level strategic planning from low-level procedural execution in human cognition. We uncover a compelling two-phase dynamic: initially, a model is constrained by procedural correctness and must improve its low-level skills. The learning bottleneck then decisively shifts, with performance gains being driven by the exploration and mastery of high-level strategic planning. This insight exposes a core inefficiency in prevailing RL algorithms like GRPO, which apply optimization pressure agnostically and dilute the learning signal across all tokens. To address this, we propose Hierarchy-Aware Credit Assignment (HICRA), an algorithm that concentrates optimization efforts on high-impact planning tokens. Our extensive experiments validate that HICRA significantly outperforms strong baselines, and offer deep insights into how reasoning advances through the lens of strategic exploration.
@@ 1 Introduction
Our investigation is guided by a key insight: RL does not train models de novo. It fine-tunes base models already imbued with priors from pre-training on vast corpora of human-written solutions. These solutions inherently encodes the hierarchical structure of human reasoning – a highly efficient cognitive strategy evolved under biological constraints. This prompts us to ask: does RL unlock advanced reasoning by (re-)discovering this hierarchical structure as a promising pathway for solving math problems?
To test this hypothesis, we analyze the RL training process through the lens of hierarchical reasoning. Drawing a parallel to the cognitive architecture of the human brain (Fig. 1), which separates high-level, deliberate strategic planning from the rapid execution of learned procedures (Murray et al., 2014; Zeraati et al., 2023; Huntenburg et al., 2018), we propose a decomposition of model-generated tokens into two functional hierarchy :
@@
Our analysis across eight text-only and vision-language models confirms this hypothesis, revealing a consistent two-phase dynamic that explains the emergence of this reasoning hierarchy in LMs. We find the optimization pressure of RL is not static; instead, its learning frontier shifts. Initially, the process is constrained by procedural correctness. A single calculation error can invalidate an entire solution, creating a powerful learning signal that compels the model to first master low-level execution tokens. Once proficiency in these foundational skills is achieved, the learning bottleneck shifts to strategic planning. We find that exploring and mastering the use of planning tokens is what unlocks significant and sustained improvements in reasoning ability.
This emergent two-phase mechanism provides a unifying framework for the puzzling phenomena observed in RL training. It explains ”aha moments” as the discovery and internalization of high-level strategic reasoning strategies, such as self-reflections. It also accounts for the ”length-scaling” effect, as employing more sophisticated strategies – involving thorough planning and logical backtracing – naturally elongates the reasoning trace with structured, strategic deliberation. Notably, it provides a unified perspective to understand the complex token entropy dynamics across different models, through the lens of high-impact planning tokens and gradually confident execution tokens.
@@
Contributions. In this work, we advance the understanding of how LLMs learn to reason via RL. We demonstrate that the learning process is not monolithic but an emergent two-phase learning dynamic driven by the hierarchical priors in base models and solution structure of the reasoning tasks. This insight reveals that the true bottleneck for advanced reasoning is the mastery of high-level strategic planning, which current agnostic credit assignment methods neglect. To bridge this gap, we pioneer with an original and simple solution, HICRA. Through extensive experiments across LLMs and VLMs, we not only validate the effectiveness of HICRA, but also offer deep insights into how HICRA works through the lens of strategic exploration.
@@ 2 The Emergent Reasoning Hierarchy
Guiding Insight: The pre-training priors and the inherent structure of reasoning tasks create a strong inductive bias. For the task of math problem-solving, hierarchical reasoning proves an efficient and prominent strategy, which is discovered through RL training and unlocks advanced reasoning.
@@ 2.1 A Functional Proxy for the Reasoning Hierarchy
To analyze this reasoning hierarchy, we must first distinguish high-level strategic planning from low-level procedural execution within the model’s generated tokens. This is challenging because a token’s function is defined by its context, not its intrinsic meaning.
To address this gap, we draw inspiration from human cognition. When a person reasons through a problem, we easily identify their strategic thinking by its function. A phrase like, “Let’s try a different approach,” functions as a high-level strategic maneuver that guides the problem-solving direction. In contrast, a phrase like, “so we add 5 to both sides,” is a low-level procedural step. Inspired by this functional distinction, we introduce Strategic Grams as a functional proxy to circumvent the difficulty of formally defining what is a “planning token”.
Strategic Grams (SGs) are defined as $n$-grams that function as a single semantic unit to guide the logical flow. We use n-grams because they capture the phrasal nature of strategic language (e.g., ”let’s consider the case”) which is lost at the single-token level. These SGs facilitate three main types of logical moves: (a) deduction, (b) branching, and (c) backtracing, as we show in Figure 2.
Given a collection of SGs, our classification heuristic is straightforward: A token is classified as a strategic planning token if it is part of a Strategic Gram in the current context. All other tokens are classified as procedural execution tokens.
For simplicity, we use “execution tokens” to encompass all non-planning tokens, including concrete calculations, formatting, and other procedural language.
A key challenge is identifying a set of SGs in a principled and automated manner, avoiding the subjectivity of manual annotation or reliance on proprietary models. Our approach is based on a key insight: SGs function as the reusable scaffolding of a reasoning process (Fig. 2). This functional role imparts a distinct statistical signature: SGs should appear frequently across a wide range of different solutions but be used sparingly within any single solution. However, linguistic diversity of strategic language presents another hurdle: a single strategic intent can be expressed in numerous ways.
Our pipeline is meticulously designed to overcome these challenges. It first groups semantically equivalent n-grams to consolidate diverse phrasing and then identifies which of these consolidated concepts exhibit the statistical signature of strategic scaffolding. Specifically, we construct the SG set via the following three-step procedure:
@@
    Semantic Clustering: We first extract all $n$-grams (where $n\in[3,5]$) from a large corpus of successful reasoning solutions. Using the sentence_transformers package, we project each $n$-gram into a semantic embedding space using a pre-trained sentence transformer, and then apply a clustering algorithm to this embedding space. This step groups lexically diverse but semantically equivalent $n$-grams into a single cluster, directly addressing the challenge of linguistic diversity.
@@
    Identification by Frequency: To identify the clusters that represent reusable reasoning patterns, we analyze their frequency at the corpus level. For each semantic cluster, we compute its Cluster Document Frequency (Cluster DF): the frequency of unique solutions that contain at least one $n$-gram from that cluster.
  3. 3.
    SG Construction: We filter for clusters with top 20% Cluster DF, implying that these SGs are common across many problems. The union of all $n$-grams within these high-frequency clusters constitutes our final set of Strategic Grams.
This data-driven procedure yields a high-precision functional proxy, not an exhaustive lexicon of SGs. To validate the robustness of this approach, we conducted a sensitivity analysis by randomly removing 30% of the identified SGs. The resulting learning dynamic curves remained qualitatively identical (see Appendix), confirming that our SG set is sufficiently representative to reveal the core learning dynamics.
@@ 2.2 Emergence of the Reasoning Hierarchy
Building on our functional proxy for reasoning, we examine the learning dynamics of RL for LLM reasoning and finds an intriguing parallel with human-like hierarchical reasoning. Our empirical analysis – conducted consistently across different model families, Qwen2.5-7B (Yang et al., 2024), Qwen3-4B (Yang et al., 2025), Llama-3.1-8B (Grattafiori et al., 2024), Qwen2.5-VL-7B (Bai et al., 2025), MiMO-VL-7B (Xiaomi, 2025) – reveals that enhanced reasoning is not a monolithic process, but driven by an evolution of the learning frontiers.
The learning process exhibit two overlapping phases: it often begins with a rapid consolidation of procedural reliability, conducive to the widespread low-level tokens. This is followed by a sustained period where the greatest potential for improvement shifts to the exploration of high-level strategic reasoning, which serves as the true engine of advanced performance.
2.2.1 Forging Reliable Low-level Skills
The initial phase of RL training is dedicated to mastering the basics. The model must first build a reliable engine for low-level skills, e.g., formatting, performing calculations and other procedural steps. To observe this, we track two key metrics on the execution tokens:
  • Relative Perplexity: Perplexity, the exponentiated average negative log-likelihood, measures model surprise. A lower value signifies higher confidence. We normalize the perplexity by its initial value to compare the rates of change in planning tokens and execution tokens.
  • Token-Level Entropy: The Shannon entropy of the policy’s next-token distribution, $H(\pi(\cdot|x_{<t}))$, measures its uncertainty. High entropy signals active exploration over the vocabulary at the next-token, while low entropy suggests confident exploitation.
@@ 2.2.1 Forging Reliable Low-level Skills
Notably, we find that this phase of low-level skill consolidation might be absent or shot in models with stronger capacity, as evident in MiMO-VL-Instruct and Qwen-4B-Instruct. This also supports the argument that the primary driver of RL is indeed the exploration of strategic planning. We refer the reader to check the full analysis of training dynamics across eight models in the appendix.
@@ 2.2.2 Steering the Skills with Strategic Planning
Once the model becomes procedurally reliable, its performance gains are primarily driven by its ability to explore and deploy a diverse set of high-level strategies. To track this shift, we analyze the planning tokens using two key metrics. We compute the Semantic Entropy of strategic grams – the Shannon Entropy of the frequency distribution of strategic grams – to quantify the diversity of the model’s high-level strategic plans (illustrated in Fig. 4). To isolate procedural variety, we compute the conditional entropy of subsequent procedural n-grams given a preceding strategic gram. This second metric shows how varied is the subsequent procedural steps for a preceding strategic move.
@@
Figure 4: Comparison of Token Entropy and Semantic Entropy. (Left) Token-level Entropy is computed over the distribution of next-token probability. (Right) Semantic Entropy is computed as the Shannon Entropy over the frequency distribution of n-grams. Intuitively, Semantic Entropy gathers tokens by their semantic function and measures the semantic diversity. Token-level entropy is not de-duplicated by semantic meanings, and is thus dominated by vast amount of high-frequency low-level tokens.
@@
Explaining Puzzling Phenomena. This emergent reasoning hierarchy provides a unified explanation for previously observed behaviors.
@@
Semantic Entropy: A Good Compass for Exploration. The trends in Figure 3 also highlight a critical flaw in using aggregate token-level entropy (column 2) to track exploration.
@@
The aggregate token-level entropy is dominated by the vast majority of low-level execution tokens. As the model become confident in procedural steps or low-level skills, the entropy of these tokens naturally decreases, pulling the global average down.
Unluckily, the decrease in token-level entropy sometimes mislead practitioners into the conception of declined exploration. This is incorrect, however, as it contradicts the fact of increasing exploration in strategic plans (semantic entropy of planning tokens) and the improving reasoning performance. Figure 4 visualize the difference between token-level entropy and semantic entropy, demonstrating that, a model can be very predictable in its next-token choice (token entropy) under a given context but still create diverse semantic structures and arguments.
In Section 4.2.3, we compare semantic entropy with token entropy and Pass@K. Our results show that semantic entropy avoids the flaws in token entropy by directly measuring diversity at the semantic level of meaningful strategic units. Its trend accurately reflects the expansion of the model’s strategic playbook, making it a more reliable diagnostic tool for tracking genuine exploration and predicting sustained performance improvements. It also complements Pass@K metric with further benefits. We refer interested readers to the appendix for a full analysis of training dynamics across eight LLM and VLMs and the deeper insights into RL training and exploration.
@@ 3 HICRA: Hierarchy-Aware Credit Assignment
Our empirical analysis reveals a fundamental insight: RL improves reasoning by rediscovering and operationalizing the strategic layer of reasoning inherited from the model’s pre-training priors. The learning process is characterized by a dynamic shift in its learning frontiers. Initially, the model is constrained by procedural correctness, but as it masters these foundational skills, the frontier for performance improvement shifts to the exploration and mastery of high-level strategic planning.
@@
Formulation. We introduce Hierarchy-Aware Credit Assignment (HICRA), an algorithm that builds upon the GRPO framework to allocate credit based on the reasoning hierarchy. In GRPO, given a query $\mathbf{q}$ from a dataset $\mathcal{D}$, the policy $\pi_{\theta}$ generates a set of $G$ output trajectories $\{\mathbf{o}_{1},\dots,\mathbf{o}_{G}\}$. The advantage for a token $o_{i,t}$ at timestep $t$ in trajectory $\mathbf{o}_{i}$ is the group-normalized reward:
$\hat{A}_{i,t}=R(\mathbf{q},\mathbf{o}_{i})-\frac{1}{G}\sum_{j=1}^{G}R(\mathbf{q},\mathbf{o}_{j})\vskip-8.5359pt$
@@
Connection to Strategic Exploration. The core mechanism of HICRA engineers more effective exploration by reshaping the policy update’s target distribution. A standard policy gradient (Williams, 2004) update nudges the policy $\pi_{\theta_{old}}$ toward an implicit target distribution $\pi^{\ast}$ defined by the advantage function described as follows (the derivation is included in the appendix):
@@
This anisotropic reshaping fosters a potent virtuous feedback loop: (a) the policy is incentivized to explore the subspace of strategic plans more thoroughly; (b) this leads to the faster discovery of effective reasoning patterns; and (c) when these strategies yield high rewards, the amplified advantage ensures they are strongly reinforced, cementing the model’s planning capabilities far more efficiently. We also validate the effects of HICRA in exploration through experiments in Section 4.